OpenStudy (anonymous):

Find side b if side a = 8.5 inches and side c = 15 inches.

OpenStudy (anonymous):



OpenStudy (nurali):

Pythagorean Theorem: a^2 + b^2 = c^2 for right triangles. So c^2 - a^2 = b^2 15^2 - 8.5^2 = b^2 152.75 = b^2 b = sqrt(152.75) b = 12.36

OpenStudy (anonymous):

thanks i am ganna posted anothr one

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

k